Find four consecutive integers such that the sum of the second and the fourth is 17 less than thrice the first.
Answer by ankor@dixie-net.com(12701)

Find four consecutive integers:
x, (x+1), (x+2), (x+3)
:
Just write what is says:
the sum of the second and the fourth is 17 less than thrice the first.
(x+1) + (x+3) = 3x - 17
x + x + 1 + 3 = 3x - 17
2x + 4 = 3x - 17
2x - 3x = -17 - 4
-x = -21
x = 21: 22, 23, 24 are the numbers
:
:
Check solution using the statement:
"the sum of the second and the fourth is 17 less than thrice the first."
22 + 24 = 3(21) - 17
46 = 63 - 17